OpenStudy (anonymous):

TES YOUR SKILLS. At a school supply store a binder cost $5 and a pencile pouches cost $2, 42 of these items are sold for total sales of $144. which system represents this how many binders were sold? A.b+p=1445b+2p=4282 bindersB.b+p=422b+5p=14422 bindersC.b-p=422b+5p=144 26 bindersD.b+p=425b+2p=14420 binders

OpenStudy (texaschic101):

5b + 2p = 144 b + p = 42 --> p = 42 - b now sub 42 - b in for p in the first equation 5b + 2(42 - b) = 144 5b + 84 - 2b = 144 3b = 144 - 84 3b = 60 b = 60/3 b = 20 ANSWER : D : b + p = 42.....5b + 2p = 144.....20 binders